Evaluate or simplify (n-1)!
Answer by Alan3354(69443)

Evaluate or simplify (n-1)!
--------------
= (n-1)*(n-2)*(n-3)*(n-4)...*1
===============
eg, 6! = 6*5*4*3*2*1 = 720
It's called factorial.